# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v2

@@ 5.1. Quality measures for causes
$\begin{array}[]{rclcl}\operatorname{\mathit{precision}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\ \Diamond\mathsf{Eff}\ |\ \Diamond\mathsf{Cause}\ )&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{tp}^{\mathfrak{S}}+\mathsf{fp}^{\mathfrak{S}}}\\[4.30554pt] \operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})&=&\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\ \Diamond\mathsf{Cause}\ |\ \Diamond\mathsf{Eff}\ )&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{tp}^{\mathfrak{S}}+\mathsf{fn}^{\mathfrak{S}}}\\[4.30554pt] \mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{\displaystyle\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl(\Diamond(\mathsf{Cause}\wedge\Diamond\mathsf{Eff})\bigr)}{\displaystyle\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl((\neg\lozenge\mathsf{Cause})\wedge\lozenge\mathsf{Eff}\bigr)}&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{fn}^{\mathfrak{S}}}.\end{array}$
Note that for these definitions we make some respective assumptions on the scheduler. We assume
@@
[…]
@@ 5.2. Computation schemes for the quality measures for fixed cause set
For this section, we assume a fixed PR cause $\mathsf{Cause}$ is given and address the problem to compute its quality values. The first observation is, that all quality measures are preserved by the switch from $\mathcal{M}$ to $\mathcal{M}_{[\mathsf{Cause}]}$ as well as the transformations of $\mathcal{M}_{[\mathsf{Cause}]}$ to an MDP that satisfies conditions (A1)-(A3) of Section 4.2. In the following Lemmata 28 and 29 we show that the quality measures $\operatorname{\mathit{recall}},\mathit{covrat}$ and $\mathit{fscore}$ of a fixed $\mathsf{Cause}$ are compatible with the model transformations from section 4. These are, on one hand a transformation to $\mathcal{M}_{[\mathsf{Cause}]}$, which only considers the minimal probability to reach $\mathsf{Eff}$ starting from $\mathsf{Cause}$, and on the other hand a transformation to an MDP $\mathcal{N}$ satisfying (A1)-(A3), which has no end components and has exactly four terminal states $\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}},\mathsf{noeff}_{\mathsf{fp}},\mathsf{noeff}_{\mathsf{tn}}$.
@@
We obtain $\operatorname{\mathit{recall}}_{\mathcal{M}}(\mathsf{Cause})\leqslant\operatorname{\mathit{recall}}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause})$ and the analogous statements for the coverage ratio and the f-score.
@@
$\begin{array}[]{rcl}\operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&\ \geqslant&\operatorname{\mathit{recall}}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause})\\ \mathit{covrat}_{\mathcal{M}}(\mathsf{Cause})&\geqslant&\mathit{covrat}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause})\end{array}$
@@
$\frac{1}{\operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})}=1+\frac{1}{\mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})}\ \ \text{and thus}\ \ \mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})=1/\left(\frac{1}{\operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})}-1\right)$
@@ 5.3. Quality-optimal probability-raising causes
where $\mathsf{fn}^{\mathfrak{S}}_{C}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl((\neg\lozenge C)\wedge\lozenge\mathsf{Eff}\bigr)$ and $\mathsf{tp}^{\mathfrak{S}}_{C}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl(\Diamond(C\wedge\Diamond\mathsf{Eff})\bigr)$. If $C$ is a cause where $\mathsf{fn}^{\mathfrak{S}}_{C}$ is positive then
$\mathit{covrat}^{\mathfrak{S}}(C)\ =\ \frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}}\quad\text{and}\quad\operatorname{\mathit{recall}}^{\mathfrak{S}}(C)\ =\ \frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}+\mathsf{tp}^{\mathfrak{S}}_{C}}$
@@
$\frac{p}{q}<\frac{p^{\prime}}{q^{\prime}}\qquad\text{iff}\qquad\frac{p}{p+q}<\frac{p^{\prime}}{p^{\prime}+q^{\prime}}$
@@
$\frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}}$ is minimal iff $\frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}+\mathsf{tp}^{\mathfrak{S}}_{C}}$ is minimal
@@
We assign weight $1$ to all states in $U=\mathsf{Eff}\cup\{\mathsf{noeff}_{\mathsf{fp}}\}$ and weight 0 to all other states of $\mathcal{K}$. Let $V=\{\mathsf{eff}_{\mathsf{cov}}\}$. Then, $f=\mathrm{E}^{\min}_{\mathcal{K}}(\boxplus V)$ and an MD-scheduler $\mathfrak{S}$ for $\mathcal{K}$ such that $\mathrm{E}^{\mathfrak{S}}_{\mathcal{K}}(\boxplus V)=f$ are computable in polynomial time. Let $\mathcal{C}_{\gamma}$ denote the set of states $c\in\mathcal{C}$ where $\mathfrak{S}(c)=\gamma$ and let $\mathsf{Cause}$ be the set of states $c\in\mathcal{C}_{\gamma}$ where $\mathcal{M}$ has a path satisfying $(\neg\mathcal{C}_{\gamma})\,\mathrm{U}\,c$. Then, $\mathsf{Cause}$ is an SPR cause of $\mathcal{M}$. With arguments as in Section 5.2 we obtain $\mathit{fscore}(\mathsf{Cause})=2/(f{+}2)$. It remains to show that $\mathsf{Cause}$ is f-score-optimal. Let $C$ be an arbitrary SPR cause. Then, $C\subseteq\mathcal{C}$. Let $\mathfrak{T}$ be the MD-scheduler for $\mathcal{K}$ that schedules $\gamma$ in $C$ and $\alpha$ for all other states of $\mathcal{K}$. Then, $\mathit{fscore}(C)=2/(f^{\mathfrak{T}}{+}2)$ where $f^{\mathfrak{T}}=\mathrm{E}^{\mathfrak{T}}_{\mathcal{K}}(\boxplus V)$. Hence, $f\leqslant f^{\mathfrak{T}}$, which yields $\mathit{fscore}(\mathsf{Cause})\geqslant\mathit{fscore}(C)$.
@@
  • $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff}\mid(\neg C)\,\mathrm{U}\,c)=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond\mathsf{Eff}\mid(\neg C)\,\mathrm{U}\,c)$ for all $c\in C$ for which the values are defined,
@@
For the direction $\Leftarrow$”, first recall that any subset of $\mathcal{C}$ satisfying (M) is an SPR cause for $\mathsf{Eff}$ in $\mathcal{N}$ by definition of $\mathcal{C}$. Now, let $\mathfrak{S}$ be a scheduler for $\mathcal{N}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})>0$. Let $\Gamma$ be the set of finite $\mathfrak{S}$-paths $\gamma$ in the sub-MDP $\mathcal{N}^{\min}_{E}$ such that $\mathfrak{S}$ chooses an action in $\mathit{Act}_{\mathcal{N}}(\mathit{last}(\gamma))\setminus\mathit{Act}^{\min}(\mathit{last}(\gamma))$ with positive probability after $\gamma$ where $\mathit{last}(\gamma)$ denotes the last state of $\gamma$. Let
$q\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}\sum_{\gamma\in\Gamma}\sum_{\alpha\in\mathit{Act}_{\mathcal{N}}(\mathit{last}(\gamma))\setminus\mathit{Act}^{\min}(\mathit{last}(\gamma))}P_{\mathcal{N}}(\gamma)\cdot\mathfrak{S}(\gamma)(\alpha).$
So, $q$ is the overall probability that a state-action pair from $\Pi$ is chosen under $\mathfrak{S}$. We now define a scheduler $\mathfrak{T}$ for $\mathcal{K}$: For each $\gamma\in\Gamma$ ending in a state $s$ and each $\alpha\in\mathit{Act}_{\mathcal{N}}(s)\setminus\mathit{Act}^{\min}(s)$, the scheduler $\mathfrak{T}$ chooses action $\beta_{(s,\alpha)}$ in $\mathsf{init}_{\mathcal{K}}$ with probability $P_{\mathcal{N}}(\gamma)\cdot\mathfrak{S}(\gamma)(\alpha)/q$. When reaching a state $t$ afterwards, $\mathfrak{T}$ behaves like $\mathit{res}(\mathfrak{S},\gamma\,\alpha\,t)$ afterwards. Note that by definition this indeed defines a probability distribution over the actions in the initial state $\mathsf{init}_{\mathcal{K}}$.
@@
The graph structure of $\mathcal{M}$ is indeed a tree and $\mathcal{M}$ can be constructed from the values $A,A_{1},\ldots,A_{n}$ and $B,B_{1},\ldots,B_{n}$ in polynomial time. Moreover, for $\mathsf{Eff}=\{\mathsf{eff}_{\mathsf{unc}}\}\cup\{\mathsf{eff}_{i}:i=0,1,\ldots,n\}$ we have:
@@
We will define polynomial-time computable values $p_{0},p_{1},\ldots,p_{n}\in\ ]0,1[$ (where $p_{i}=P(\mathsf{init},s_{i})$), $w_{1},\ldots,w_{n}\in\ ]0,1[$ (where $w_{i}=P(s_{i},\mathsf{eff}_{i})$) and auxiliary variables $\delta\in\ ]0,1[$ and $\lambda>1$ such that:
@@ 6.1.1. Checking causality and existence of reachability PR causes
To explore how this change of definition influences the previously established results for GPR and SPR causes, we have to clarify how effects will be represented. We use deterministic Rabin automata (DRAs) as they are expressive enough to capture all $\omega$-regular languages and they are deterministic, which will allow us to form well-behaved products of the automata with MDPs. Let $\mathcal{M}$ be an MDP, $\mathsf{rEff}$ an effect given by the DRA $\mathcal{A}_{\mathsf{rEff}}$ and $\mathsf{Cause}\subseteq S$ a cause candidate. As a special case we again have Markov chains with no non-deterministic choices. Then, the conditions (rG) and (rS) can easily be checked by computing the corresponding probabilities in polynomial time (see [BKKM14] for algorithms to compute conditional probabilities in MCs for path properties.) We now consider the case where non-deterministic choices exist. We will provide a model transformation of $\mathcal{M}$ using the DRA such that the resulting MDP has no end components and the effect is a reachability property again.
@@
Now, let $\mathfrak{T}$ be a scheduler for $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$. The scheduler $\mathfrak{T}$ can be mimicked by a scheduler $\mathfrak{S}$ for $\mathcal{M}$: As long as $\mathfrak{T}$ moves through the MEC-quotient of $\mathcal{N}^{\prime}$, the scheduler $\mathfrak{S}$ follows this behavior by leaving MECs through the corresponding actions in $\mathcal{M}$. Whenever $\mathfrak{T}$ moves to one of the states $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{eff}_{\mathsf{unc}}$, the last step begins in a state $s_{\mathcal{E}}$ obtained from a MEC $\mathcal{E}$ of $\mathcal{M}$. In this case, $\mathfrak{S}$ will stay in $\mathcal{E}$ while ensuring with probability $1$ that the resulting run is accepted by $\mathcal{A}_{\mathsf{rEff}}$. Similarly, if $\mathfrak{T}$ moves to $\mathsf{noeff}_{\mathsf{tn}}$ or $\mathsf{noeff}_{\mathsf{fp}}$, the scheduler $\mathfrak{S}$ for $\mathcal{M}$ stays in the corresponding MEC and realizes the acceptance condition of $\mathcal{A}_{\mathsf{rEff}}$ with probability 0. Vice versa, a scheduler $\mathfrak{S}$ for $\mathcal{M}$ can be mimicked by a scheduler $\mathfrak{T}$ for $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ analogously. So in an end component $\mathcal{E}$, if $\mathfrak{S}$ stays in $\mathcal{E}$ and ensures that the resulting run is accepted by $\mathcal{A}_{\mathsf{rEff}}$ with probability $p_{1}$, stays in $\mathcal{E}$ and ensures that the resulting run is not accepted by $\mathcal{A}_{\mathsf{rEff}}$ with probability $p_{2}$, and leaves $\mathcal{E}$ with probability $p_{3}$, $\mathfrak{T}$ will move to the corresponding state $\mathsf{eff}_{\mathsf{cov}}$ or $\mathsf{eff}_{\mathsf{unc}}$ with probability $p_{1}$, to $\mathsf{noeff}_{\mathsf{tn}}$ or $\mathsf{noeff}_{\mathsf{fp}}$ with probability $p_{2}$, and leave take actions leaving $s_{\mathcal{E}}$ to other states with the same probability distribution with which $\mathfrak{S}$ takes the leaving actions of $\mathcal{E}$.

### v6

@@ 5.1. Quality measures for causes
$\begin{array}[]{rclcl}\operatorname{\mathit{precision}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\ \Diamond\mathsf{Eff}\ |\ \Diamond\mathsf{Cause}\ )&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{tp}^{\mathfrak{S}}+\mathsf{fp}^{\mathfrak{S}}}\\[8.61108pt] \operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})&=&\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\ \Diamond\mathsf{Cause}\ |\ \Diamond\mathsf{Eff}\ )&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{tp}^{\mathfrak{S}}+\mathsf{fn}^{\mathfrak{S}}}\\[8.61108pt] \mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{\displaystyle\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl(\Diamond(\mathsf{Cause}\wedge\Diamond\mathsf{Eff})\bigr)}{\displaystyle\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl((\neg\lozenge\mathsf{Cause})\wedge\lozenge\mathsf{Eff}\bigr)}&=&\frac{\displaystyle\mathsf{tp}^{\mathfrak{S}}}{\displaystyle\mathsf{fn}^{\mathfrak{S}}}.\end{array}$
Note that for these quality measures we make some respective assumptions on the scheduler as we assume
@@
$\mathit{covrat}(\mathsf{Cause})\ =\ \inf_{\mathfrak{S}}\mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause}),\quad\mathit{fscore}(\mathsf{Cause})\ =\ \inf_{\mathfrak{S}}\mathit{fscore}^{\mathfrak{S}}(\mathsf{Cause}).$
@@ 5.2. Computation schemes for the quality measures for fixed cause set
For this section, we assume a fixed PR cause $\mathsf{Cause}$ is given and address the problem to compute its quality values. The first observation is, that all quality measures are preserved by the switch from $\mathcal{M}$ to $\mathcal{M}_{[\mathsf{Cause}]}$ as well as the transformations of $\mathcal{M}_{[\mathsf{Cause}]}$ to an MDP that satisfies conditions (A1)-(A3) of Section 4.2. In the following Lemmata 28 and 29 we show that the quality measures $\operatorname{\mathit{recall}},\mathit{covrat}$ and $\mathit{fscore}$ of a fixed $\mathsf{Cause}$ are compatible with the model transformations from Section 4. These are, on one hand a transformation to $\mathcal{M}_{[\mathsf{Cause}]}$, which only considers the minimal probability to reach $\mathsf{Eff}$ starting from $\mathsf{Cause}$, and on the other hand a transformation to an MDP $\mathcal{N}$ satisfying (A1)-(A3), which has no end components and has exactly four terminal states $\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}},\mathsf{noeff}_{\mathsf{fp}},\mathsf{noeff}_{\mathsf{tn}}$. These four terminal states exactly correspond to the four entries of the confusion matrix (Figure 13).
@@
We obtain the claimed inequalities, e.g. $\operatorname{\mathit{recall}}_{\mathcal{M}}(\mathsf{Cause})\leqslant\operatorname{\mathit{recall}}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause})$.
@@
$\operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})\ \geqslant\ \operatorname{\mathit{recall}}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause}),$ and $\mathit{covrat}_{\mathcal{M}}(\mathsf{Cause})\ \geqslant\ \mathit{covrat}_{\mathcal{M}_{[\mathsf{Cause}]}}(\mathsf{Cause})$
@@
$\frac{1}{\operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})}=1+\frac{1}{\mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})}\ \text{thus}\ \mathit{covrat}^{\mathfrak{S}}(\mathsf{Cause})=1/\left(\frac{1}{\operatorname{\mathit{recall}}^{\mathfrak{S}}(\mathsf{Cause})}-1\right).$
@@ 5.3. Quality-optimal probability-raising causes
where $\mathsf{fn}^{\mathfrak{S}}_{C}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl((\neg\lozenge C)\wedge\lozenge\mathsf{Eff}\bigr)$ and $\mathsf{tp}^{\mathfrak{S}}_{C}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}\bigl(\Diamond(C\wedge\Diamond\mathsf{Eff})\bigr)$. If $C$ is a cause where $\mathsf{fn}^{\mathfrak{S}}_{C}$ is positive then $\mathit{covrat}^{\mathfrak{S}}(C)\ =\ \frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}}$ and $\operatorname{\mathit{recall}}^{\mathfrak{S}}(C)\ =\ \frac{\mathsf{tp}^{\mathfrak{S}}_{C}}{\mathsf{fn}^{\mathfrak{S}}_{C}+\mathsf{tp}^{\mathfrak{S}}_{C}}$.
@@
$\frac{p}{q}<\frac{p^{\prime}}{q^{\prime}}\qquad\text{iff}\qquad\frac{p}{p+q}<\frac{p^{\prime}}{p^{\prime}+q^{\prime}}.$
@@
[…]
@@
We assign weight $1$ to all states in $U=\mathsf{Eff}\cup\{\mathsf{noeff}_{\mathsf{fp}}\}$ and weight 0 to all other states of $\mathcal{K}$. Let $V=\{\mathsf{eff}_{\mathsf{cov}}\}$. Then, $f=\mathrm{E}^{\min}_{\mathcal{K}}(\boxplus V)$ and an MD-scheduler $\mathfrak{S}$ for $\mathcal{K}$ such that $\mathrm{E}^{\mathfrak{S}}_{\mathcal{K}}(\boxplus V)=f$ are computable in polynomial time. Let $\mathcal{C}_{\gamma}$ denote the set of states $c\in\mathcal{C}$ where $\mathfrak{S}(c)=\gamma$ and let $\mathsf{Cause}$ be the set of states $c\in\mathcal{C}_{\gamma}$ where $\mathcal{M}$ has a path satisfying $(\neg\mathcal{C}_{\gamma})\,\mathrm{U}\,c$. Then, $\mathsf{Cause}$ is an SPR cause of $\mathcal{M}$. With arguments as in Section 5.2 we obtain $\mathit{fscore}(\mathsf{Cause})=2/(f{+}2)$.
It remains to show that $\mathsf{Cause}$ is f-score-optimal. Let $C$ be an arbitrary SPR cause. Then, $C\subseteq\mathcal{C}$. Let $\mathfrak{T}$ be the MD-scheduler for $\mathcal{K}$ that schedules $\gamma$ in $C$ and $\alpha$ for all other states of $\mathcal{K}$. Then, $\mathit{fscore}(C)=2/(f^{\mathfrak{T}}{+}2)$ where $f^{\mathfrak{T}}=\mathrm{E}^{\mathfrak{T}}_{\mathcal{K}}(\boxplus V)$. Hence, $f\leqslant f^{\mathfrak{T}}$, which yields $\mathit{fscore}(\mathsf{Cause})\geqslant\mathit{fscore}(C)$.
@@
  • $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff}\mid(\neg C)\,\mathrm{U}\,c)=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond\mathsf{Eff}\mid(\neg C)\,\mathrm{U}\,c)$ for all $c\in C$ where the values are defined,
@@
For the direction $\Leftarrow$”, first recall that any subset of $\mathcal{C}$ satisfying (M) is an SPR cause for $\mathsf{Eff}$ in $\mathcal{N}$ by definition of $\mathcal{C}$. Now, let $\mathfrak{S}$ be a scheduler for $\mathcal{N}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})>0$. Let $\Gamma$ be the set of finite $\mathfrak{S}$-paths $\gamma$ in the sub-MDP $\mathcal{N}^{\min}_{E}$ such that $\mathfrak{S}$ chooses an action in $\mathcal{A}=\mathit{Act}_{\mathcal{N}}(\mathit{last}(\gamma))\setminus\mathit{Act}^{\min}(\mathit{last}(\gamma))$ with positive probability after $\gamma$ where $\mathit{last}(\gamma)$ denotes the last state of $\gamma$. Let
$q\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}\sum_{\gamma\in\Gamma}\sum_{\alpha\in\mathcal{A}}P_{\mathcal{N}}(\gamma)\cdot\mathfrak{S}(\gamma)(\alpha).$
So, $q$ is the overall probability that a state-action pair from $\Pi$ is chosen under $\mathfrak{S}$. We now define a scheduler $\mathfrak{T}$ for $\mathcal{K}$: For each action $\gamma\in\Gamma$ ending in a state $s$ and each action $\alpha\in\mathit{Act}_{\mathcal{N}}(s)\setminus\mathit{Act}^{\min}(s)$, the scheduler $\mathfrak{T}$ chooses action $\beta_{(s,\alpha)}$ in $\mathsf{init}_{\mathcal{K}}$ with probability $P_{\mathcal{N}}(\gamma)\cdot\mathfrak{S}(\gamma)(\alpha)/q$. When reaching a state $t$ afterwards, $\mathfrak{T}$ behaves like $\mathit{res}(\mathfrak{S},\gamma\,\alpha\,t)$ afterwards. Note that by definition this indeed defines a probability distribution over the actions in the initial state $\mathsf{init}_{\mathcal{K}}$.
@@
The graph structure of $\mathcal{M}$ is indeed a tree and $\mathcal{M}$ can be constructed from the values $A,A_{1},\ldots,A_{n}$ and $B,B_{1},\ldots,B_{n}$ in polynomial time. For $\mathsf{Eff}=\{\mathsf{eff}_{\mathsf{unc}}\}\cup\{\mathsf{eff}_{i}:i=0,1,\ldots,n\}$ we have:
@@
Next, we will consider the polynomial-time computable values $p_{0},p_{1},\ldots,p_{n}\in\ ]0,1[$ (where $p_{i}=P(\mathsf{init},s_{i})$), $w_{1},\ldots,w_{n}\in\ ]0,1[$ (where $w_{i}=P(s_{i},\mathsf{eff}_{i})$) and auxiliary variables $\delta\in\ ]0,1[$ and $\lambda>1$ such that:
@@ 6.1.1. Checking causality and existence of reachability PR causes
To explore how this change of definition influences the previously established results for GPR and SPR causes, we have to clarify how effects will be represented. We use deterministic Rabin automata (DRAs) as they are expressive enough to capture all $\omega$-regular languages and they are deterministic, which will allow us to form well-behaved products of the automata with MDPs. Let $\mathcal{M}$ be an MDP, $\mathsf{rEff}$ an effect given by the DRA $\mathcal{A}_{\mathsf{rEff}}$ and $\mathsf{Cause}\subseteq S$ a cause candidate. As a special case we again have Markov chains with no non-deterministic choices. Then, the conditions (rG) and (rS) can easily be checked by computing the corresponding probabilities in polynomial time (see [BKKM14] for algorithms to compute conditional probabilities in MCs for path properties). We now consider the case where non-deterministic choices exist. We will provide a model transformation of $\mathcal{M}$ using the DRA such that the resulting MDP has no end components and the effect is a reachability property again.
@@
Now, let $\mathfrak{T}$ be a scheduler for $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$. The scheduler $\mathfrak{T}$ can be mimicked by a scheduler $\mathfrak{S}$ for $\mathcal{M}$: As long as $\mathfrak{T}$ moves through the MEC-quotient of $\mathcal{N}^{\prime}$, the scheduler $\mathfrak{S}$ follows this behavior by leaving MECs through the corresponding actions in $\mathcal{M}$. Whenever $\mathfrak{T}$ moves to one of the states $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{eff}_{\mathsf{unc}}$, the last step begins in a state $s_{\mathcal{E}}$ obtained from a MEC $\mathcal{E}$ of $\mathcal{M}$. In this case, $\mathfrak{S}$ will stay in $\mathcal{E}$ while ensuring with probability $1$ that the resulting run is accepted by $\mathcal{A}_{\mathsf{rEff}}$. Similarly, if $\mathfrak{T}$ moves to $\mathsf{noeff}_{\mathsf{tn}}$ or $\mathsf{noeff}_{\mathsf{fp}}$, the scheduler $\mathfrak{S}$ for $\mathcal{M}$ stays in the corresponding MEC and realizes the acceptance condition of $\mathcal{A}_{\mathsf{rEff}}$ with probability 0. Vice versa, a scheduler $\mathfrak{S}$ for $\mathcal{M}$ can be mimicked by a scheduler $\mathfrak{T}$ for $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ analogously. So in an end component $\mathcal{E}$, if $\mathfrak{S}$ stays in $\mathcal{E}$ and ensures that the resulting run is accepted by $\mathcal{A}_{\mathsf{rEff}}$ with probability $p_{1}$, stays in $\mathcal{E}$ and ensures that the resulting run is not accepted by $\mathcal{A}_{\mathsf{rEff}}$ with probability $p_{2}$, and leaves $\mathcal{E}$ with probability $p_{3}$, $\mathfrak{T}$ will move to the corresponding state $\mathsf{eff}_{\mathsf{cov}}$ or $\mathsf{eff}_{\mathsf{unc}}$ with probability $p_{1}$, to $\mathsf{noeff}_{\mathsf{tn}}$ or $\mathsf{noeff}_{\mathsf{fp}}$ with probability $p_{2}$, and take actions leaving $s_{\mathcal{E}}$ to other states with the same probability distribution with which $\mathfrak{S}$ takes the leaving actions of $\mathcal{E}$.